# Phenomenology of laminar acoustic streaming jets

### v1

@@ II.1 Definition of the streaming problem
We consider a steady and axisymmetric Eckart streaming flow of velocity field $\bm{u_{s}}$ in a closed cylindrical cavity of diameter $D_{c}$ and length $L_{c}$ (Fig. 1). The fluid is Newtonian of density $\rho$ and kinematic viscosity $\nu$.
@@
A plane circular transducer of diameter $D_{s}$ is placed at $x=0$ and radiates a beam-shaped sound field centred on the cylinder axis. The sound waves, of power $P_{ac}$ and of wavelength $\lambda$, propagate through the medium at phase speed $c$. In addition to the decay caused by diffraction, the amplitude of the acoustic pressure waves drops exponentially along the axis due to viscous and thermal effects occurring within the bulk of the fluid. This decay rate is given by the sound attenuation coefficient [36, 31]:
@@
where $\eta$ and $\mu=\rho\nu$ are the bulk and dynamic viscosities, respectively. As formerly derived by Lighthill [1], the theoretical axial momentum flux of an Eckart streaming jet flowing in a semi-infinite domain increases with the distance $x_{d}$ from the transducer, and asymptotically reaches its maximum $M_{th,max}=P_{ac}/c$ at infinite $x_{d}$. In the case where the effective beam length matches $L_{c}$, only a fraction $\left(1-e^{-2\alpha L_{c}}\right)$ of $M_{th,max}$ is actually injected into the fluid volume. In the present work, we focus on a $L_{c}>\alpha$ case to maximize the jet momentum flux. Using $\left(\alpha,\lambda,M_{th,max},\rho,\nu,D_{s},D_{c},L_{c},\bm{u_{s}}\right)$ as design parameters, and $D_{s}$, $\rho D_{s}^{3}$, $D_{s}^{2}/\nu$ as length, mass and time scales, respectively, we describe the acoustic propagation and hydrodynamics problems using six dimensionless groups [31]:
@@
The last term in Eq. (3b) is the acoustic force driving the flow. The spatial structure of the forcing is set by the normalized acoustic intensity $\bm{\widetilde{I}}$ dictated by the acoustic problem (see Sec. II.2). The magnitude of the forcing, relative to the magnitude of the viscous forces, is given by the acoustic Grashof number $Gr_{ac}$ [23]:
@@
Finally, the values adopted for the dimensionless parameters (2) are listed in Table 1. The dimensionless acoustic attenuation coefficient $N/L=\alpha D_{s}$ and the diffraction parameter $S$ are chosen to match the experimental setup in Ref. [31], so that ${Gr_{ac}\boldsymbol{\widetilde{I}}}$ is representative of the forcing field of former experiments. An important aspect of our work is the moderately large value of $N$ for which $\alpha^{-1}\lesssim L_{c}$, i.e., the beam is entirely attenuated before reaching the end of the fluid domain, yielding forcing-free regions at far distances from the source. Typically, at $x=L$, the remaining acoustic power is about 0.01 % of the power radiated by the transducer. We finally choose $D=6L$ to minimize the effects of the lateral wall on the jet.
@@ II.2 Equations governing the spatial distribution of the attenuated acoustic force field
    Sound waves are irrotational, i.e., ×$\bm{\nabla}\times\bm{u_{ac}}=\bm{0}$. This is a common assumption in acoustics [36, 37, 39]; it holds in the bulk of the fluid but breaks down in a thin layer near solid boundaries, including at the source. There, a viscous layer of dimensionless thickness $\delta=\sqrt{\nu\lambda/\left(\pi cD_{s}^{2}\right)}$ forces the acoustic velocity fluctuation $\bm{u_{ac}}$ to match the boundary velocity. This is typically the case near the vibrating membrane of the transducer radiating the acoustic waves. For the experiment in Ref. [31], on which we base our forcing field structure, $\delta\approx 1.3\times 10^{-5}$. As $\delta$ is several orders of magnitudes smaller than the cavity and flow length scales, the supplementary attenuation and possible streaming flow within that layer can be safely neglected [40].
@@
    Nonlinear effects (e.g., wave steepening) are discarded. The waves are thus linear, and the superposition principle applies.
@@
    Sound attenuation is due to the viscosity of the medium only. Asymptotic analysis [39] shows that thermally-induced attenuation is proportional to $\left(\gamma-1\right)$, with $\gamma$ the adiabatic index. For common liquids such as water, $\gamma\approx 1$, hence thermally-induced attenuation can be dropped [9] (this approximation however does not hold for liquid metals). We also discard any relaxation-based attenuation, e.g., through the excitation of the vibrational and rotational modes of the fluid molecules [36].
@@
$\bm{\widetilde{I}}$ =$=\Re\left\{\widetilde{p_{ac}}\,\bm{\widetilde{u}_{ac}}^{*}\right\}\,,$ (5a)
[…]
@@
The resulting acoustic intensity field exhibits strong gradients for $x\leq L_{F}$ (see Fig. 11 (b) and Moudjed et al. [42]), with $L_{F}=1.22/\left(4S\right)$ the Fresnel length locating the transition between the near ($x<L_{F}$) and far ($x>L_{F}$) acoustic fields. Since (i) the beam and source radii are equal to 0.5 at $x=0$ and (ii) the beam expands radially at a rate $S$ due to diffraction [31], an approximate description of the beam radius variation along $x$ is:
@@
The acoustic intensity $\widetilde{I}_{x}(x,r=0)$ along the beam axis is decaying mainly due to this radial expansion effect but also from attenuation. The drop of $P_{ac}$ along the beam due to attenuation yields an exponential contribution to the decay of $\widetilde{I}_{x}(x,r=0)$. From Eq. (36), the length scale for this exponential decay is defined by $L_{\alpha}=L/\left(2N\right)\approx 166.7$: at $x=L_{\alpha}$, $\|\bm{\widetilde{I}}\|$ has dropped by 63 % compared to the unattenuated case.
@@ III The role of the acoustic force attenuation
In the literature, the acoustic attenuation has frequently been discarded in the model of the acoustic force to derive analytical 1D-1C Eckart streaming velocity profiles within a weak attenuation assumption [2, 3, 35, 47, 48, 21, 40, 49]. We want to emphasize the importance of the contribution of the attenuation in the spatial structure of the acoustic force, even for cases where the jet observation length is small compared to $L_{\alpha}$. For this purpose we will focus on numerical calculations of the on-axis velocity profile, the radius of the jet and the momentum flux of the jet, by computing Eqs. (5) considering two cases: $N/L=0$ in Eq. (5d) (without attenuation, as it is often considered in the literature) and $N/L\neq 0$ (with attenuation, as it will be considered from Sec. IV onwards). We shall then compare the numerical results to the theoretical predictions of Lighthill [1] regarding the jet momentum flux.
@@
[…]
@@ IV Characterization of long-range acoustic streaming jets
The flow is steady, axisymmetric and laminar. The evolution of the jet velocity along its axis is characterized by a strong acceleration near the transducer (yellow and red regions in the bottom plot of Fig. 5). The irregularities witnessed in the on-axis velocity profile for $1\leq x\leq 10$ are caused by the large longitudinal gradients of the force in the acoustic near field (Fig. 11 (b)) and were discussed in previous works [27, 42]. On the axis, the jet velocity then reaches a peak value, whose position $x_{p}$ does not coincide with either of the acoustic quantities $L_{F}$ or $L_{\alpha}$. At $x_{p}$, the Reynolds number is maximum and is equal to 150.9 for $Gr_{ac}=5\times 10^{3}$. Downstream $x_{p}$, the on-axis jet velocity decays smoothly as the distance from the transducer increases (blue region in the bottom plot of Fig. 5). Despite flowing over large distances compared to $L_{\alpha}$ (i.e., up to distances over which the sound beam has been nearly entirely attenuated), the jet impinges the downstream wall with a velocity of nearly 10 % of its peak value. The jet impingement occurring at $x=L$ then gives rise to a large vortex of weak velocity amplitude. In the following, we shall only consider the jet itself and each of the flow features identified on the axial jet velocity profile, from the fluid region close to the transducer up to the downstream wall.
@@ IV.1 Inertia-dominated region of accelerating fluid
For $0.5\leq x\leq x_{p}$, former studies have shown that the observed fluid acceleration results from a balance between inertia effects and the acoustic force [31, 50]. In this previous work, the authors considered that since the length of the acceleration region is small compared to $L_{\alpha}$, it was legitimate to neglect both the attenuation and the beam diffraction. The order of magnitude $U_{x,i}$ of the inviscid on-axis velocity was then given by [31]:
@@
An order of magnitude accounting for viscous forces can be obtained from the first component of the Navier-Stokes equation, Eq. (3b). Because of the large cavity diameter $D$, pressure forces are negligible, hence:
@@
Since the scaling law for $U_{x,v}$ is sought for $x\ll L_{\alpha}$, and the beam enlargement is small in the considered region, the spatial variations of $\widetilde{I}_{x}$ caused by attenuation and diffraction can be neglected in Eq. (10). Besides, Fig. 3 (b) shows that $R_{jet}\sim R_{beam}\sim 1/2$ near the transducer. Further taking $\frac{\partial\left(U_{x,v}^{2}\right)}{\partial x}\sim U_{x,v}^{2}/x$, Eq. (9) then reduces to a quadratic polynomial of positive root
@@ IV.2 Scaling of the on-axis velocity peak position and amplitude
The region of jet acceleration is followed by a velocity peak $u_{x,p}$ at $x_{p}$. Both $x_{p}$ and $u_{x,p}$ display a weak dependency on $Gr_{ac}$ that we shall now investigate. At $x=x_{p}$ and $r=0$, both $\partial u_{x}/\partial x$ and $\partial u_{x}/\partial r$ are zero, causing the left-hand side of the steady form of the Navier-Stokes equation (3b) to vanish. In other words, $u_{x,p}$ is controlled by a local balance between the acoustic and viscous forces:
@@
Figure 7: Evolution of the on-axis velocity peak position $x_{p}$ and amplitude $u_{x,p}$ with $Gr_{ac}$. (a) Comparison between the peak positions observed in the numerical simulations (black filled symbols), and the predictions when neglecting viscous forces ($x_{p,i}$, blue line, Eq. (15a)) and considering them ($x_{p,v}$, red line, Eq. (16)) in the region upstream $x_{p}$. As the velocity peak lies in the far-acoustic field, both $x_{p,i}$ and $x_{p,v}$ are determined using the far-field approximation (36) for the on-axis intensity $\widetilde{I}_{x}$. The quantity represented on the $x$-axis is chosen to assess the validity of Eq. (15a). (b) Comparison between the peak amplitudes observed in the numerical simulations (black filled symbols), the experiments from Dentry et al. [30] (empty symbols) and the theoretical predictions (solid lines) when viscous forces upstream $x_{p}$ are neglected ($U_{x,p_{i}}$, blue, Eq. (15b)) and taken into account ($U_{x,p_{v}}$, red, Eqs. (16) and (14)). The quantity plotted on the $x$-axis is chosen to assess the validity of Eq. (15b). In (b), the data of Dentry et al. [30] are plotted considering an equivalent height $H_{s}$ of the source varying with the frequency. The parameters $Gr_{ac}$ and $S$ are thus based on $H_{s}$ for the experimental points (instead of $D_{s}$ for the numerical simulations).
@@
$U_{x,p_{i}}\sim\frac{1}{2}\left[\frac{1}{64}\left(\frac{1.22\pi}{8\,S}\right)^{4}\right]^{0.1}Gr_{ac}^{0.6}\,.$ (15b)
@@
The obtained prediction $x_{p_{v}}$, plotted as a red line in Fig. 7 (a), appears as a better estimate of the peak position: the difference between $x_{p}$ and $x_{p_{v}}$ remains between 11 % and 13 % for $Gr_{ac}^{0.2}/S^{0.8}\geq 105$. At lower $Gr_{ac}$, that discrepancy is however larger. The reason is that $R_{jet}$ varies slightly with $Gr_{ac}$. For decreasing values of $Gr_{ac}$, the jet velocity also decreases. As a consequence, the effect of viscous diffusion on the radial enlargement of the jet becomes more important: at $L_{F}$, $R_{jet}$ measured for $Gr_{ac}=10^{3}$ is 64 % larger than for $Gr_{ac}=5\times 10^{4}$ (not shown). Thus, considering the same $R_{jet}$ for all $Gr_{ac}$ in Eq. (9) leads to an overestimation of the viscous forces at low $Gr_{ac}$. Using a larger value of $R_{jet}$ in Eq. (9) would also shift the range of $Gr_{ac}$ values over which this equation is valid towards lower values.
The same loss of accuracy at low $Gr_{ac}$ is found for $U_{x,p_{v}}$, albeit to a lesser extent (Fig. 7 (b)). Although $U_{x,p_{i}}$ provides an acceptable discrepancy of 8.7 % with the velocity peak magnitudes observed in the numerical simulations, its variation with $Gr_{ac}$ is further improved when viscous forces are taken into account upstream the peak. The difference between $U_{x,p_{v}}$ and $u_{x,p}$ indeed falls below 3 % for $Gr_{ac}^{0.6}/S^{0.4}\geq 4\times 10^{2}$. Values issued from recent experimental observations [30] are also plotted as empty symbols. They were obtained in the case of non-axisymmetric jet generated using Rayleigh Surface Acoustic Waves devices operating at different frequencies. The equivalent acoustic source is rectangular, with a height $H_{s}$ depending on the frequency; $Gr_{ac}$ and $S$ are thus based on $H_{s}$ for plotting the experimental points. These experimental data clearly display a similar dependence of $u_{x,p}$ on $Gr_{ac}$, and thus provide an experimental evidence that the velocity peak amplitude indeed satisfies a local balance between viscous and acoustic forces.

### v2

@@ II.1 Definition of the streaming problem
We consider a steady and axisymmetric Eckart streaming flow of dimensional velocity field $\bm{u_{s}}$ in a closed cylindrical cavity of diameter $D_{c}$ and length $L_{c}$ (Fig. 1). The fluid is Newtonian of density $\rho$ and kinematic viscosity $\nu$.
@@
A plane circular transducer of diameter $D_{s}$ is placed at $x=0$ and radiates a beam-shaped sound field centered on the cylinder axis. The sound waves, of power $P_{ac}$ and of wavelength $\lambda$, propagate through the medium at a phase speed $c$. In addition to the decay caused by diffraction, the amplitude of the acoustic pressure waves drops exponentially along the axis due to viscous and thermal effects occurring within the bulk of the fluid. This decay rate is given by the sound attenuation coefficient [36, 31]:
@@
where $\eta$ and $\mu=\rho\nu$ are the bulk and dynamic viscosities, respectively, $\gamma$ is the adiabatic index and $Pr=\mu c_{p}/\kappa$ is the Prandtl number with $c_{p}$ the specific heat capacity at constant pressure and $\kappa$ the thermal conductivity. As formerly derived by Lighthill [1], the theoretical axial momentum flux of an Eckart streaming jet flowing in a semi-infinite domain increases with the distance $x_{d}$ from the transducer, and asymptotically reaches its maximum $M_{th,max}=P_{ac}/c$ at infinite $x_{d}$. In the case where the effective beam length matches $L_{c}$, only a fraction $\left(1-e^{-2\alpha L_{c}}\right)$ of $M_{th,max}$ is actually injected into the fluid volume. In the present work, we focus on a $L_{c}>\left(1/\alpha\right)$ case to maximize the jet momentum flux. Using $\left(\alpha,\lambda,M_{th,max},\rho,\nu,D_{s},D_{c},L_{c},\bm{u_{s}}\right)$ as dimensional design parameters, and $D_{s}$, $\rho D_{s}^{3}$, $D_{s}^{2}/\nu$ as length, mass and time scales, respectively, we describe the acoustic propagation and hydrodynamics problems using six dimensionless groups [31]:
@@
The last term in Eq. (3b) is the acoustic force driving the flow  and is valid far from solid boundaries and for any acoustic wave shape [9]. The spatial structure of the forcing is set by the normalized acoustic intensity $\bm{\widetilde{I}}$ dictated by the acoustic problem (see Sec. II.2). $\bm{\widetilde{I}}$ is normalized by the maximum unattenuated plane wave intensity $I_{max}$ on the transducer axis (Eq. (34)), giving $\|\bm{\widetilde{I}}\|\leq 1$ everywhere in the fluid domain. The magnitude of the forcing, relative to the magnitude of the viscous forces, is then given by the acoustic Grashof number $Gr_{ac}$ [23]:
@@
The values of the dimensionless parameters (2) are listed in Table 1. We choose the same diffraction parameter $S=0.03$ and attenuation coefficient $N/L=3\times 10^{-3}$ as in the water experiments of Moudjed et al. [31]. Similar parameters values are found in the experiments of Refs. [27, 29, 38], in which $S\in\left[0.014,0.036\right]$ and $N/L\in\left[5.9\times 10^{-3},18.8\times 10^{-3}\right]$. The acoustic forcing structure in our simulations is thus representative of those commonly found in streaming jets experiments.
An important aspect of our work is the moderately large $N$. In contrast to Refs. [27, 29, 38, 31], we choose $N>1$. The beam is thus entirely attenuated before reaching the end of the fluid domain, yielding forcing-free regions at far distances form the source. Typically, at $x=L$, the remaining acoustic power is about 0.01 % of the power radiated by the transducer. To minimize the effects of the lateral wall on the jet, we choose $D\approx L/3$.
Finally, we take $10^{3}\leq Gr_{ac}\leq 5\times 10^{4}$, which fits in the ranges explored by Kamakura et al. [27], Mitome [29], Frenkel et al. [38], Moudjed et al. [31]. Some of these authors also investigated larger $Gr_{ac}$. We limited ourselves to $Gr_{ac}\leq 5\times 10^{4}$, as computing steady flows at greater $Gr_{ac}$ in such a large cavity becomes prohibitively expensive.
@@ II.2 Equations governing the spatial distribution of the attenuated acoustic force field
    Sound waves are irrotational, i.e., $\bm{\nabla}\times\bm{u_{ac}}=\bm{0}$. This is a common assumption in acoustics [36, 37, 40]; it holds in the bulk of the fluid but breaks down in a thin layer near solid boundaries, including at the source. There, a viscous layer of dimensionless thickness $\delta=\sqrt{\nu\lambda/\left(\pi cD_{s}^{2}\right)}$ forces the acoustic velocity fluctuation $\bm{u_{ac}}$ to match the boundary velocity. This is typically the case near the vibrating membrane of the transducer radiating the acoustic waves. For the experiment in Ref. [31], on which we base our forcing field structure, $\delta\approx 1.3\times 10^{-5}$. As $\delta$ is several orders of magnitudes smaller than the cavity and flow length scales, the supplementary attenuation and possible streaming flow within that layer can be safely neglected [41].
@@
    We assume that $L_{\alpha}$ is smaller than the acoustic shock formation distance: the sound waves are thus linear and the superposition principle applies. This modelling choice removes the need for defining a frequency [31] and acoustic nonlinearity parameters [40], thus making our scaling study as generic as possible.
@@
    Sound attenuation is due to the viscosity of the medium only. Asymptotic analysis [40] shows that thermally-induced attenuation is proportional to $\left(\gamma-1\right)$. For common liquids such as water, $\gamma\approx 1$, hence thermally-induced attenuation can be dropped [9] (this approximation however does not hold for liquid metals). We also discard any relaxation-based attenuation, e.g., through the excitation of the vibrational and rotational modes of the fluid molecules [36].
@@
$\bm{\widetilde{I}}$ $=\Re\left\{\widetilde{p_{ac}}\,\bm{\widetilde{u}_{ac}}^{*}\right\}\,$, (5a)
$\widetilde{p}_{ac}$ $=j\frac{k^{2}S}{8\times 1.22\pi^{2}}\int_{0}^{1/2}\int_{0}^{2\pi}\frac{e^{-jk\|\bm{d}\|}}{\|\bm{d}\|}r^{\prime}\mathrm{d}r^{\prime}\mathrm{d}\theta^{\prime}\,$, (5b)
$\bm{\widetilde{u}_{ac}}$ $=\frac{1}{4\pi}\int_{0}^{1/2}\int_{0}^{2\pi}\left(1+jk\|\bm{d}\|\right)\frac{e^{-jk\|\bm{d}\|}}{\|\bm{d}\|^{2}}\frac{\bm{d}}{\|\bm{d}\|}r^{\prime}\mathrm{d}r^{\prime}\mathrm{d}\theta^{\prime}\,$, (5c)
@@
The resulting acoustic intensity field exhibits strong gradients for $x\leq L_{F}$ (see Fig. 11 (b) and Moudjed et al. [43]), with $L_{F}=1.22/\left(4S\right)\approx 10$ the Fresnel length locating the transition between the near ($x<L_{F}$) and far ($x>L_{F}$) acoustic fields. Since (i) the beam and source radii are equal to 0.5 at $x=0$ and (ii) the beam expands radially at a rate $S$ due to diffraction [31], an approximate description of the beam radius variation along $x$ is:
@@
The acoustic intensity $\widetilde{I}_{x}(x,r=0)$ along the beam axis is decaying mainly due to this radial expansion effect but also from attenuation. The drop of $P_{ac}$ along the beam due to attenuation yields an exponential contribution to the decay of $\widetilde{I}_{x}(x,r=0)$. From Eq. (36), the length scale for this exponential decay is thus defined by $L_{\alpha}=L/\left(2N\right)\approx 166.7$. The ratio $L_{F}/L_{\alpha}\approx 0.06$ is thus marginal, meaning that the near acoustic field structure is barely affected by attenuation. In our work, we choose to focus on the scaling of the jet velocity with $Gr_{ac}$ and thus keep $L_{F}/L_{\alpha}$ constant. For more details on the effect of $L_{F}/L_{\alpha}$ on the jet velocity near the source, the interested reader shall refer to Daru et al. [33].
@@ III The role of the acoustic force attenuation
In the literature, the acoustic attenuation has frequently been discarded in the model of the acoustic force to derive analytical 1D-1C (one-dimensional, one-component) Eckart streaming velocity profiles within a weak attenuation assumption [2, 3, 35, 48, 49, 21, 41, 50]. We want to emphasize the importance of the contribution of the attenuation in the spatial structure of the acoustic force, even for cases where the jet observation length is small compared to $L_{\alpha}$. For this purpose we will focus on numerical calculations of the on-axis velocity profile, the radius of the jet and the momentum flux of the jet, by computing Eqs. (5) considering two cases: $N/L=0$ in Eq. (5d) (without attenuation, as it is often considered in the literature) and $N/L\neq 0$ (with attenuation, as it will be considered from Sec. IV onwards). We shall then compare the numerical results to the theoretical predictions of Lighthill [1] regarding the jet momentum flux.
@@
From these numerical results, we can now study the evolution along $x$ of the axial jet momentum flux defined as $\int_{0}^{2\pi}\int_{0}^{D/2}u_{x}^{2}rdrd\theta$. As illustrated in Fig. 4 for $Gr_{ac}=10^{3}$, when $N/L\neq 0$, the axial jet momentum flux (filled symbols) saturates for $x\geq 500$ up to $x=1300$. On the contrary, setting $N/L=0$ leads to a linear increase with $x$ of the axial momentum flux (empty symbols). This is in agreement with the theoretical prediction by Lighthill [1] for an axisymmetric Eckart streaming jet driven by an attenuated sound beam in a semi-infinite medium: the evolution with $x$ of the axial momentum flux, then defined as $\int_{0}^{2\pi}\int_{0}^{\infty}u_{x}^{2}rdrd\theta$, is equal to the integral of the acoustic force along the beam if the pressure gradient is neglected. This momentum injected by the acoustic forcing in the fluid will be denoted $M_{th}$, and obeys to the following expression:
@@ IV Characterization of long-range acoustic streaming jets
The flow is steady, axisymmetric and laminar. The evolution of the jet velocity along its axis is characterized by a strong acceleration near the transducer (yellow and red regions in the bottom plot of Fig. 5). The irregularities witnessed in the on-axis velocity profile for $1\leq x\leq 10$ are caused by the large longitudinal gradients of the force in the acoustic near field (Fig. 11 (b)) and were discussed in previous works [27, 43]. On the axis, the jet velocity then reaches a peak value, whose position $x_{p}$ does not coincide with either of the acoustic quantities $L_{F}$ or $L_{\alpha}$. At $x_{p}$, the Reynolds number is maximum and is equal to 150.9 for $Gr_{ac}=5\times 10^{3}$. Downstream $x_{p}$, the on-axis jet velocity decays smoothly as the distance from the transducer increases (blue region in the bottom plot of Fig. 5). Despite flowing over large distances compared to $L_{\alpha}$ (i.e., up to distances over which the sound beam has been nearly entirely attenuated), the jet impinges the downstream wall with a velocity of nearly 10 % of its peak value. The jet impingement occurring at $x=L$ then gives rise to a large vortex of weak velocity amplitude. In the following, we shall only consider the jet itself and each of the flow features identified on the axial jet velocity profile, from the fluid region close to the transducer up to the downstream wall. We shall rely on an order-of-magnitude approach [51] (i) to determine the approximate magnitude of the on-axis velocity $u_{x}$ in the jet regions identified in the bottom plots of Fig. 5, (ii) to explain the evolution of $u_{x}$ in these regions based on a force balance, and (iii) to determine how $u_{x}$ scales with the parameters of the problem.
@@ IV.1 Inertia-dominated region of accelerating fluid
For $0.5\leq x\leq x_{p}$, former studies have shown that the jet accelerates due to a balance between the acoustic and inertia forces [31, 52]. In these previous works, the authors considered that since the length of the acceleration region is small compared to $L_{\alpha}$, it was legitimate to neglect both the attenuation and the beam diffraction [33]. The order of magnitude $U_{x,i}$ of the inviscid on-axis velocity was then given by [31]:
@@
An order of magnitude accounting for viscous forces can be obtained from the first component of the Navier-Stokes equation, Eq. (3b). On $0.5\leq x\leq x_{p}$, the jet is almost purely axial and the order of magnitude of the inertia term in Eq. (3b) is given by $u_{x}\partial u_{x}/\partial x$ [31]. Furthermore, the pressure forces are negligible due to the large cavity diameter $D$. From an order of magnitude standpoint, the Navier-Stokes equations (3b) thus give:
@@
Since both diffraction and attenuation are weak over the jet region of interest (i.e., $S\ll 1$ and $x\ll L_{\alpha}$ respectively), the spatial variations of $\widetilde{I}_{x}$ caused by attenuation and diffraction can be neglected in Eq. (10). Besides, Fig. 3 (b) shows that $R_{jet}\sim R_{beam}\sim 1/2$ near the transducer. Further taking $\frac{\partial\left(U_{x,v}^{2}\right)}{\partial x}\sim U_{x,v}^{2}/x$, Eq. (9) then reduces to a quadratic polynomial of positive root
@@ IV.2 Scaling of the on-axis velocity peak position and amplitude
The region of jet acceleration is followed by a velocity peak $u_{x,p}$ at $x_{p}$. Both $x_{p}$ and $u_{x,p}$ display a weak dependency on $Gr_{ac}$ that we shall now investigate. At $x=x_{p}$ and $r=0$, both $\partial u_{x}/\partial x$ and $\partial u_{x}/\partial r$ are zero, causing the left-hand side of the steady form of the Navier-Stokes equation (3b) to vanish. For any $Gr_{ac}$, $u_{x,p}$ is thus controlled by a local balance between the acoustic and viscous forces:
@@
Figure 7: Evolution of the on-axis velocity peak position $x_{p}$ and amplitude $u_{x,p}$ with $Gr_{ac}$. (a) Comparison between the peak positions observed in the numerical simulations (black filled symbols), and the predictions when neglecting viscous forces ($x_{p,i}$, blue line, Eq. (15a)) and considering them ($x_{p,v}$, red line, Eq. (16)) in the region upstream $x_{p}$. As the velocity peak lies in the far-acoustic field, both $x_{p,i}$ and $x_{p,v}$ are determined using the far-field approximation (36) for the on-axis intensity $\widetilde{I}_{x}$. The quantity represented on the $x$-axis is chosen to assess the validity of Eq. (15a). (b) Comparison between the peak amplitudes observed in the numerical simulations (black filled symbols), the experiments from Dentry et al. [30] (empty symbols) and the theoretical predictions (solid lines) when viscous forces upstream $x_{p}$ are neglected ($U_{x,p_{i}}$, blue, Eq. (15b)) and taken into account ($U_{x,p_{v}}$, red, Eqs. (16) and (14)). The quantity plotted on the $x$-axis is chosen to assess the validity of Eq. (15b). In (b), the data of Dentry et al. [30] correspond to a jet of rectangular cross section. The beam driving the jet is radiated by a rectangular Surface Acoustic Waves (SAW) device of height $H_{s}$ that varies with frequency. For the data points of Dentry et al. [30], $Gr_{ac}$ and $S$ are thus based on $H_{s}$ (instead of $D_{s}$ in our numerical simulations).
@@
$U_{x,p_{i}}\sim\frac{1}{2}\left[\frac{1}{64}\left(\frac{1.22\pi}{8\,S}\right)^{4}\right]^{\frac{1}{10}}Gr_{ac}^{\frac{3}{5}}\,.$ (15b)
@@
The prediction $x_{p_{v}}$ (red curve in Fig. 7 (a)) yields a better estimate of the peak position for $Gr_{ac}^{\frac{1}{5}}/S^{\frac{4}{5}}\geq 105$ (i.e., $Gr_{ac}\geq 10^{4}$). At lower $Gr_{ac}$, the larger discrepancy is caused by the dependency of $R_{jet}$ on $Gr_{ac}$. We recall that $x_{p_{v}}$ relies on the velocity scaling law (11) obtained by assuming that $R_{jet}\approx 0.5$ i.e., the jet width is set by the beam width for all $Gr_{ac}$. While this approximation holds for $Gr_{ac}\geq 10^{4}$, we however observed that the increased viscous diffusion at lower $Gr_{ac}$ widens the jet. For instance, $R_{jet}$ measured at $x=L_{F}$ for $Gr_{ac}=10^{3}$ is 64 % larger than for $Gr_{ac}=10^{4}$ (not shown), resulting in a 128 % difference on the viscous force estimate. Since the estimate of the velocity peak position is obtained by balancing the acoustic with the viscous forces, the error on $R_{jet}$ thus causes $x_{p_{v}}$ to diverge away from the numerical simulation results for $Gr_{ac}<10^{4}$. Thus, considering the same $R_{jet}$ for all $Gr_{ac}$ in Eq. (9) leads to an overestimation of the viscous forces at low $Gr_{ac}$.
The same loss of accuracy at low $Gr_{ac}$ is found for $U_{x,p_{v}}$, albeit to a lesser extent (Fig. 7 (b)). Although $U_{x,p_{i}}$ provides an acceptable discrepancy of 8.7 % with the velocity peak magnitudes observed in the numerical simulations, its variation with $Gr_{ac}$ is further improved when viscous forces are taken into account upstream the peak. The difference between $U_{x,p_{v}}$ and $u_{x,p}$ indeed falls below 3 % for $Gr_{ac}^{3/5}/S^{2/5}\geq 4\times 10^{2}$. For comparison, we added the experimental data points of Dentry et al. [30] (empty symbols) to Fig. 7 (b). In their experiments, the authors used rectangular Rayleigh Surface Acoustic Waves (SAW) devices to drive streaming jets. Contrary to our numerical setup, the jets in Ref. [30] were forced by a beam of rectangular cross section and were therefore non-axisymmetric. These differences in the forcing and flow geometries inevitably affect the magnitudes of the forces to which the jet is subjected, causing the values of $u_{x,p}$ to differ. Nevertheless, the data points of Ref. [30] scale similarly with $Gr_{ac}$ and $S$: this proves that $u_{x,p}$ indeed satisfies a local balance between the acoustic and viscous forces, as assumed by the scaling laws (15b) and (14). Improving the match between these scaling laws and the $u_{x,p}$ of Ref. [30] would require developments specific to their peculiar jet geometry that are out of the scope of the present work.